OpenStudy (anonymous):

Deck of Cards • 4 Suits: • Each suit has 13 cards: A, 2 , 3, 4, 5, 6, 7, 8, 9, 10, J, Q, K • 52 cards • P(Jack) = 4/52 = 1/13 • P(Heart) = 13/52 = 1/4 • P(Jack AND Heart) =

OpenStudy (anonymous):

y is the answer 1/52

OpenStudy (anonymous):

Are these cards from the same deck, meaning you draw first either a jack or a heart and then the other, or do you draw them from different decks at the same time?

OpenStudy (anonymous):

iono it doesnt specify

OpenStudy (anonymous):

alright I assume that it means from the same deck so here it goes. To take the probability of to seperate events occuring, you multiply their probabilities together. The first event is either drawing a jack or an ace. this probability is given by:

OpenStudy (anonymous):

1/13 + 1/13 = 0.15 The second event occurs when you have already gotten rid of one of the 52 cards, so now you have a 51 card deck. If you pulled an ace the first time there will be 4 jacks and vice versa. so the probability of pulling the second card will be 4/51 = 0.078 then to get the overall probability you multiply them both together Prob = (2/13)(4/51) Prob = 0.012 or 1.2 % chance

OpenStudy (zarkon):

Jack AND Heart this means the jack of hearts. There is only 1 jack of hearts and there are 52 cards...thus the prob is 1/52

OpenStudy (anonymous):

my bad i misread the problem

OpenStudy (anonymous):

=(